# research-article-2col

### v2

@@ 1 Introduction
Sodium ions (Na+) plays a critical role in the human body and are invariably linked to the maintenance of ionic homeostasis as well as many physiological and electrochemical processes of metabolism (1, 2). For this reason, the non-invasive detection of these ions from the nuclear magnetic resonance (NMR) signal of the sodium isotope New A(spin $\frac{\text{3}}{\text{2}}$) has become an important modality for the study of metabolic regulation and biochemical status throughout the body (1, 2). In brain, New AMRI has been used to study neurodegenerative disease (1, 2, 3, 4, 5, 6, 7), tumor pathology (1, 2, 8, 9) and neurological events, such as stroke (2, 10) and traumatic brain injury (11). The intracellular and extracellular spaces in brain tissues both represent motion-restricted environments that give rise to signal contributions coming from the central and satellite transitions, which are strongly influenced by the quadrupolar interactions of the New Anuclear spin system with its surroundings (2). These dynamics result in a biexponential transverse relaxation decay, i.e. a long component ($\text{T}_{\text{2,long}}^{*}$) and a short component ($\text{T}_{\text{2,short}}^{*}$), in both the intracellular and extracellular spaces. As a consequence, both gray matter (GM) and white matter (WM) in brain will also exhibit an overall New Abiexponential transverse relaxation, as a weighted average of the relaxation processes from the intracellular and extracellular spaces. Similarly, New Alongitudinal relaxation follows the same biexponential pattern from the intracellular and extracellular spaces. However, in soft biological tissues, both the short and long components are often very close to each other and longitudinal relaxation is usually measured as a monoexponential process ($\text{T}_{\text{1,short}}\sim\text{T}_{\text{1,long}}\sim\text{T}_{\text{1}}$) in GM and WM. In fluids such as cerebrospinal fluid (CSF), the quadrupolar interaction averages to zero as a result of rapid motion allowing the signal dynamics to be often modeled as a monoexponential relaxation for both transverse and longitudinal magnetization components. Collectively, the spin ${\frac{\text{3}}{\text{2}}}$ nature of the New Anucleus and the low concentration of Na+ ions in brain tissue (on the order of 40-50 mM on average) combined with the inhomogeneous structure of the brain, makes it difficult to simultaneously quantify New Arelaxation times and density (1, 2, 12).
@@
Proton magnetic resonance fingerprinting (1H MRF) has become a popular technique for the simultaneous quantification of physical properties within a system (13, 14, 15, 16, 17). Recently, sodium MRF (New AMRF) studies in the brain have demonstrated promising initial results. Kratzer et al. (18, 19) implemented a version of New AMRF capable of quantifying relaxation parameters in CSF and brain tissue (combined GM and WM) that utilized a 3D radial sequence with variable repetition times (TR), echo times (TE) and flip angles (FA). Our group previously introduced a multipulse approach to multicompartmental Na+ concentration quantification (20), which we now expanded for quantifying New Arelaxation in the brain.
@@ 2 Material and Methods
$\frac{d}{dt}\rho(t)=-i\big{[}H,\rho(t)\big{]}-\hat{\Gamma}\{\rho(t)-\rho^{th}\},\vspace{4pt}$ (1)
where $\rho^{th}$ is the density operator of the spin system at thermal equilibrium, $H$ is the total spin Hamiltonian and $\hat{\Gamma}$ is the Redfield relaxation superoperator. The total Hamiltonian $H$ is the sum of the main Hamiltonians acting on the density operator, such as the Zeeman Hamiltonian $H_{Z}$ (interaction of the spins with the constant $B_{0}$ field), the residual quadrupolar coupling Hamiltonian $H_{Q}$ (interaction of the quadrupole moment of the nuclei with the residual average electric field gradient in anisotropic environments), and the RF field Hamiltonian $H_{1}(t)$ (interaction of the spins with time-varying transmit RF field $B_{1}^{+}(t)$). In soft quasi-isotropic brain tissues, $H_{Q}$ can be considered negligible. Since we are also operating in the rotating frame for the spin dynamics simulation, the main Hamiltonians acting on the spin system are $H_{Z}$ related to $B_{0}$ inhomogeneities only, and $H_{1}(t)$.
@@
$\begin{split}\hat{\Gamma}\{\rho-\rho^{th}\}=A\sum^{2}_{m=-2}(-1)^{m}[T_{2,m},[T_{2,-m},\rho-\rho^{th}]]\\ \times\big{(}J(m\omega)-iK(m\omega)\big{)},\end{split}\vspace{-4pt}$ (2)
where $A$ is a constant that depends on the convention used to describe spectral densities, $T_{2,m}$ is the ISTO of rank 2 and order m, $J(m\omega)$ is the spectral density function, and $K(m\omega)$ is the imaginary term associated with dynamic frequency shift, which in our case can be omitted due to its negligible observable effect in soft tissues (2). The matrix formulation of $\hat{\Gamma}$ and relaxation rates $R_{i}=\frac{1}{T_{i}}$ with $i$ = (1,short), (1,long), (2,short), (2,long), are given by (12):
@@
Figure 1: Pulse sequence diagram for 3D New AMRF. The overall pulse scheme is shown in (A). The diagram in (B) represents the variable flip angle (FA) and phase angle (PA) MRF train. There were a total of $N$ = 23 non-selective rectangular RF pulses of duration $\tau_{RF}=0.8$ ms. Each pulse was followed by a time period $\tau_{i}$. On the RF channel in (A), the gray rectangles correspond to a single FA = $\theta_{i}$ and PA = $\phi_{i}$ combination in train (B). In (B), the dark gray rectangles represent a 3-pulse inversion composite block (25) used to increase $\text{T}_{\text{1}}$ sensitivity of the sequence and improve RF homogeneity for the magnetization inversion. Within the composite block, $\tau_{1}$ = $\tau_{2}$ = 7.5 ms. The light gray rectangles indicate the 20-pulse variable FA and PA train. All FA = $\theta_{i}$ and PA = $\phi_{i}$ are listed at the top of each rectangle and $\tau_{i}$ = 15 ms for $i\geq 3$. During each delay $\tau_{i}$, a time period TE = 0.2 ms was followed by the ADC event, indicated by a yellow block on the ADC channel. A 3D spiral encoding scheme using the FLORET trajectory (20, 24, 26) (3 hubs at 45, 100 interleaves/hub), indicated in blue, was played over the gradient channels. Immediately after the ADC, a rewind gradient was also played out and then the next RF pulse in the MRF train was initiated. After $N$ pulses and $N\tau_{i}$ delays have played out, a 5-ms spoiling gradient at 70% maximum gradient strength, indicated by an orange trapezoid, was applied in all directions. The entirety of this scheme represented one TR of the sequence.
@@
Figure 1(A) presents the New AMRF sequence. System excitation was driven by a series of $N$ non-selective rectangular RF pulses with FA $\theta_{i}$ and PA $\phi_{i}$ followed by a delay $\tau_{i}$ ($i=1,2,...,N$), forming the MRF pulse train shown in Figure 1(B). An initial magnetization inversion using a 90–180–90 (dark gray) composite 3-pulse block (25) was used to increase $\text{T}_{\text{1}}$ sensitivity and improve RF homogeneity for the inversion before initiating the 20-pulse variable FA/PA train (light gray).
The RF pulse durations were fixed at $\tau_{RF}$ = 0.8 ms, and interpulse delay periods were fixed at $\tau_{i}$ = 7.5 ms within the composite block ($i=1,2$) and $\tau_{i}$ = 15 ms for the next pulses ($i\geq 3$). The interpulse delays were set according to our prior work on simultaneous 1H/23Na MRI (27, 28) and in anticipation of integrating this method into simultaneous 1H/23Na MRF.
The FLORET trajectory (20, 24, 26) was used to readout the signal followed by a rewinder to balance the gradient moment. After $N$ pulses and $N\tau_{i}$ delays, a spoiling gradient (duration = 5 ms, 70% of maximum gradient strength) was applied simultaneously in all directions to ensure complete dephasing of residual transverse magnetization prior to beginning the next TR. A delay of $\sim$300 ms was inserted between RF pulse trains to allow recovery of the longitudinal magnetization and reduce specific absorption rate (SAR). A constant TE = 0.2 ms was used throughout the sequence.
@@
Simulation of the fingerprint dictionary was performed in MATLAB R2020b on a Cray CS500-1211 cluster with Intel Xeon Gold 6148 high memory CPUs at the NYU Langone High Performance Computing Core facility (New York City, NY, USA). The simulation code is freely available in Matlab File Exchange (see Data Availability section for the link). Signals were simulated starting from thermal equilibrium and propagated under the optimized 23-pulse New AMRF train (Figure 1(B)). Parameter ranges ([begin:step:end]) to build the dictionary were $\text{T}_{\text{1,long}}$ = [20:2:74] ms, $\text{T}_{\text{1,short}}$ = [20:2:74] ms, $\text{T}_{\text{2,long}}^{*}$ = [10:2:66] ms, $\text{T}_{\text{2,short}}^{*}$ = [0.5:0.5:2.0, 2:2:66] ms, $\Delta\text{B}_{\text{1}}^{+}$ factor = [0.7:0.1:1.3] (applied as a multiplying factor to the RF amplitude FA) and $\Delta\text{f}_{\text{0}}$ = [-60:10:60] Hz. For $\text{T}_{\text{1}}$ quantification, we originally assumed that T1,long = T1,short. To insure that spectral densities where $J(0)\neq J(\omega)$ were included in the dictionary, a constant (±$\pm\Delta\text{T}_{\text{1}}$) that added 1 ms to every entry for $\text{T}_{\text{1,long}}$ and subtracted 1 ms to every entry for $\text{T}_{\text{1,short}}$ was included in the simulations.
Parameter combinations where $\text{T}_{\text{2,long}}^{*}>\text{T}_{\text{1}}$ and $\text{T}_{\text{2,short}}^{*}>\text{T}_{\text{2,long}}^{*}$ were omitted from the computation. Due to the memory expense incurred by the simulation, the dictionary was generated in two parts and concatenated. In total, 831,512 entries were generated. The total simulation time was 6.4 days.
@@
All experiments were performed at 7 T (MAGNETOM, Siemens, Erlangen, Germany) using a 16-channel transmit/receive 1H/23Na RF brain coil constructed in-house (8 1H channels, 8 New Achannels) (31).
@@
Our test phantom was constructed using a 2.3-L cylinder (outer diameter OD = 20 cm, length = 35 cm) filled with a solution of 70 mM NaCl and which contained seven 50-mL polypropylene cylinders (OD = 30 mm, length = 115 mm) arranged as 6 outer samples each containing a different concentration of NaCl and agar circling the 7th sample placed in the middle and containing a solution of 140 mM NaCl. A diagram of the phantom with sodium/agar concentrations in the axial orientation is shown in Figure 2. The regions-of-interest (ROI) in the 7 samples were generated from a 3D mask of equal diameter to each sample such that all ROI volumes would be the same.
For New AMRF, one scan consisting of 16 averages with TR = 511 ms, isotropic resolution = 5 mm and isotropic FOV = 320 mm, was acquired with the FLORET parameters: 3 hubs/45 with 100 interleaves/hub, TE = 0.2 ms, total scan time = 40:52 min.
@@
All images were reconstructed offline in MATLAB. For each channel, raw k-space data acquired during the New AMRF sequence was filtered with a Hamming kernel then reconstructed using gridding (32, 33) to produce a set of 23 complex images with a final nominal resolution 2.5$\times$2.5$\times$2.5 mm3 (128$\times$128$\times$128 matrix). The images from the 8 channels were combined using coil sensitivity profiles as described by Bydder et al. (34). The average time for image reconstruction in the brain across all five volunteers was 42 s, and 38 s in the phantom. For the brain images, an additional denoising step was performed on the complex images using the Marchenko-Pastur method (35, 36, 37). This added another 15 s to the brain image reconstruction time. Supplementary Figure S1 shows 23 axial images for a center slice in the phantom. Supplementary Figures S2 and S3 show 23 reconstructed axial images for a center slice in the brain of volunteer 5.
@@
For $\text{T}_{\text{2}}^{*}$, we applied the biexponential kernel shown in Equation 8 as descibed by Ridley et. al (39, 38).
@@
For each volunteer, images from the 1H MPRAGE DICOM datasets were co-registered to the New AMRF data using SPM12 (UCL, London, UK) (41, 28). The New AMRF image used for co-registration was the one acquired just after the first pulse in the New AMRF train. Tissue probability maps were generated from the normalized and co-registered MPRAGE images and segmented into CSF, GM and WM tissue classes using SPM12 (41, 28). The segmentated regions for CSF, GM and WM were then binarized with a 90% threshold to generate tissue ROI masks in MATLAB. These images are presented in Figure 3. The high threshold was chosen to reduce the likelihood of contamination between different tissues.
@@
Because of the dictionary size, and due to the low SNR of the sodium images, it was therefore possible that more than one match could generate a high correlation for a single voxel. To account for this, we included matches for a subset of the top correlations for each voxel $v$, and generated the final maps by calculating the correlation-weighted parameter $X_{v}$ from the dictionary of values $X_{v,i}$ corresponding to the matching correlation coefficient $w_{v,i}$ according to:
@@
SD was calculated as the mean absolute value of signal per correlation for each voxel $v$, normalized by the highest voxel intensity in the image. Weighting was then done according to Equation 9, where $X_{v,i}$ was the unweighted SD and $X_{v}$ was the weighted SD.

### v3

@@ 1 Introduction
Sodium ions (Na+) plays a critical role in the human body and are invariably linked to the maintenance of ionic homeostasis as well as many physiological and electrochemical processes of metabolism (1, 2). For this reason, the non-invasive detection of these ions from the nuclear magnetic resonance (NMR) signal of the sodium isotope New A(spin $\frac{\text{3}}{\text{2}}$) has become an important modality for the study of cellular ionic homeostasis and volume regulation, and biochemical status throughout the body (1, 2). In brain, New AMRI has been used to study neurodegenerative disease (1, 2, 3, 4, 5, 6, 7), tumor pathology (1, 2, 8, 9) and neurological events, such as stroke (2, 10) and traumatic brain injury (11).
@@
The intracellular and extracellular spaces in brain tissues both represent motion-restricted environments that give rise to signal contributions coming from the central and satellite transitions, which are strongly influenced by the quadrupolar interactions of the New Anuclear spin system with its surroundings (2). The central and satellites transitions correspond to the transitions (i.e. single quantum coherences [SQC], which can be detected as NMR signal) between the four energy levels of the spin S = 3/2 of New Aplaced in a large magnetic field, and which are associated with spin states ms = 3/2, 1/2, -1/2 and -3/2, from lowest to highest energy levels, respectively (2, 12). Transitions (3/2$\leftrightarrow$(1/2) and (-1/2$\leftrightarrow$(-3/2) are called “satellite” and transition (1/2$\leftrightarrow$(-1/2) is called “central”. These transitions occur with different relaxation times, the shorter ones corresponding to the satellite transitions in solids, or in biological tissues with motional restriction for the sodium ions (such as intracellular and extracellular spaces). Since a New ANMR or MRI experiment detect all SQCs related to these transitions, the overall New Asignal acquired in most tissues will exhibit a biexponential transverse relaxation decay, i.e. a long component ($\text{T}_{\text{2,long}}^{*}$) and a short component ($\text{T}_{\text{2,short}}^{*}$), originating in both the intracellular and extracellular spaces. As a consequence, both gray matter (GM) and white matter (WM) in brain will also exhibit an overall New Abiexponential transverse relaxation, as a weighted average of the relaxation processes from the intracellular and extracellular spaces. Similarly, New Alongitudinal relaxation follows the same biexponential pattern from the intracellular and extracellular spaces. However, in soft biological tissues, both the short and long components are often very close to each other and longitudinal relaxation is usually measured as a monoexponential process ($\text{T}_{\text{1,short}}\sim\text{T}_{\text{1,long}}\sim\text{T}_{\text{1}}$) in GM and WM(2). In fluids such as cerebrospinal fluid (CSF), the quadrupolar interaction averages to zero as a result of rapid motion allowing the signal dynamics to be often modeled as a monoexponential relaxation for both transverse and longitudinal magnetization components. Collectively, the spin ${\frac{\text{3}}{\text{2}}}$ nature of the New Anucleus and the low concentration of Na+ ions in brain tissue (on the order of 40-50 mM on average) combined with the inhomogeneous structure of the brain, makes it difficult to simultaneously quantify New Arelaxation times and density (1, 2, 13).
@@
Proton magnetic resonance fingerprinting (1H MRF) has become a popular technique for the simultaneous quantification of physical properties within a system (14, 15, 16, 17, 18). Recently, sodium MRF (New AMRF) studies in the brain have demonstrated promising initial results. Kratzer et al. (19, 20) implemented a version of New AMRF capable of quantifying relaxation parameters in CSF and brain tissue (combined GM and WM) that utilized a 3D radial sequence with variable repetition times (TR), echo times (TE) and flip angles (FA). Our group previously introduced a multipulse approach to multicompartmental Na+ concentration quantification (21), which we now expanded for quantifying New Arelaxation in the brain.
@@ 2 Material and Methods
$\frac{d}{dt}\rho(t)=-i\big{[}H,\rho(t)\big{]}-\hat{\Gamma}\{\rho(t)-\rho^{eq}\},\vspace{4pt}$ (1)
where $\rho^{eq}$ is the density operator of the spin system at thermal equilibrium, $H$ is the total spin Hamiltonian and $\hat{\Gamma}$ is the Redfield relaxation superoperator. The total Hamiltonian $H$ is the sum of the main Hamiltonians acting on the density operator, such as the Zeeman Hamiltonian $H_{Z}$ (interaction of the spins with the constant $B_{0}$ field), the residual quadrupolar interaction Hamiltonian $H_{Q}$ (interaction of the quadrupole moment of the nuclei with the residual average electric field gradient in anisotropic environments), and the RF field Hamiltonian $H_{1}(t)$ (interaction of the spins with time-varying transmit RF field $B_{1}^{+}(t)$) (2). According to recent studies (26, 27), this residual $H_{Q}$ can actually have an effect on $B_{1}^{+}$ efficiency and therefore on the results from quantitative sodium imaging. Since we will include $\Delta B_{1}^{+}$ inhomogeneities in our spin dynamics simulation to generate the fingerprint dictionary, we can assume that this anisotropic residual $H_{Q}$ is negligible in soft quasi-isotropic brain tissues in our model (2). Since we are also operating in the rotating frame for the spin dynamics simulation, the main Hamiltonians acting on the spin system are $H_{Z}$ related to $B_{0}$ inhomogeneities only, and $H_{1}(t)$.
@@
$\begin{split}\hat{\Gamma}\{\rho-\rho^{eq}\}=A\sum^{2}_{m=-2}(-1)^{m}[T_{2,m},[T_{2,-m},\rho-\rho^{eq}]]\\ \times\big{(}J(m\omega)-iK(m\omega)\big{)},\end{split}\vspace{-4pt}$ (2)
where $A$ is a constant that depends on the convention used to describe spectral densities, $T_{2,m}$ is the ISTO of rank 2 and order m, $J(m\omega)$ is the spectral density function, and $K(m\omega)$ is the imaginary term associated with dynamic frequency shift, which in our case can be omitted due to its negligible observable effect in soft tissues (2). The matrix formulation of $\hat{\Gamma}$ and relaxation rates $R_{i}=\frac{1}{T_{i}}$ with $i$ = (1,short), (1,long), (2,short), (2,long), are given by (13, 2):
@@
Figure 1: Pulse sequence diagram for 3D New AMRF. The overall pulse scheme is shown in (A). The diagram in (B) represents the variable flip angle (FA) and phase angle (PA) MRF train. There were a total of $N$ = 23 non-selective rectangular RF pulses of duration $\tau_{RF}=0.8$ ms. Each pulse was followed by a time period $\tau_{i}$. On the RF channel in (A), the gray rectangles correspond to a single FA = $\theta_{i}$ and PA = $\phi_{i}$ combination in train (B). In (B), the dark gray rectangles represent a 3-pulse inversion composite block (28) used to increase $\text{T}_{\text{1}}$ sensitivity of the sequence and improve RF homogeneity for the magnetization inversion. Within the composite block, $\tau_{1}$ = $\tau_{2}$ = 7.5 ms. The light gray rectangles indicate the 20-pulse variable FA and PA train. All FA = $\theta_{i}$ and PA = $\phi_{i}$ are listed at the top of each rectangle and $\tau_{i}$ = 15 ms for $i\geq 3$. During each time delay $\tau_{i}$, a time period TE = 0.2 ms was followed by the ADC event, indicated by a yellow block on the ADC channel. A 3D spiral encoding scheme using the FLORET trajectory (21, 25, 29) (3 hubs at 45, 100 interleaves/hub), indicated in blue, was played over the gradient channels. Immediately after the ADC, a rewind gradient was also played out and then the next RF pulse in the MRF train was initiated. After $N$ pulses and $N$ delays of duration $\tau_{i}$ have played out, a 5-ms spoiling gradient at 70% maximum gradient strength, indicated by an orange trapezoid, was applied in all directions. The entirety of this scheme represented one TR of the sequence.
@@
Figure 1(A) presents the New AMRF sequence. System excitation was driven by a series of $N$ non-selective rectangular RF pulses with FA $\theta_{i}$ and PA $\phi_{i}$ followed by a delay time $\tau_{i}$ ($i=1,2,...,N$), forming the MRF pulse train shown in Figure 1(B). An initial magnetization inversion using a 90–180–90 (dark gray) composite 3-pulse block (28) was used to increase $\text{T}_{\text{1}}$ sensitivity and improve RF homogeneity for the inversion before initiating the 20-pulse variable FA/PA train (light gray).
The RF pulse durations were fixed at $\tau_{RF}$ = 0.8 ms, and interpulse delay periods were fixed at $\tau_{i}$ = 7.5 ms within the composite block ($i=1,2$) and $\tau_{i}$ = 15 ms for the next pulses ($i\geq 3$). The interpulse delays were set according to our prior work on simultaneous 1H/23Na MRI (30, 31) and in anticipation of integrating this method into simultaneous 1H/23Na MRF.
The FLORET trajectory (21, 25, 29) was used to readout the signal followed by a rewinder to balance the gradient moment. After $N$ pulses and $N\tau_{i}$ delays, a spoiling gradient (duration = 5 ms, 70% of maximum gradient strength) was applied simultaneously in all directions to ensure complete dephasing of residual transverse magnetization prior to beginning the next TR. A delay of $\sim$300 ms was inserted between RF pulse trains to allow recovery of the longitudinal magnetization and reduce specific absorption rate (SAR). A constant TE = 0.2 ms was used throughout the sequence.
The variable 20-pulse FA/PA train was optimized using a genetic algorithm (GA) implemented in the Global Optimization Toolbox in MATLAB R2020b (The MathWorks Inc., Natick, Massachusetts, USA) on an Apple MacBook Pro (2019) laptop with a 2.4 GHz 8-Core Intel Core i9 processor. The GA operates by evolving a population of candidate solutions through the processes of selection, recombination and mutation or crossover. First, a random population is initialized and each individual is evaluated using the objective function. These individuals are scored and assigned a fitness value which is then converted into an expectation value. Next, parents for the following population are determined based on their expectation values. At each step, the GA generates a new population of offsprings from the pool of current parents. The GA terminates when the user-defined end criteria are met.
@@
Simulation of the fingerprint dictionary was performed in MATLAB R2020b on a Cray CS500-1211 cluster with Intel Xeon Gold 6148 high memory CPUs at the NYU Langone High Performance Computing Core facility (New York City, NY, USA). The simulation code is freely available in Matlab File Exchange (see Data Availability section for the link). Signals were simulated starting from thermal equilibrium and propagated under the optimized 23-pulse New AMRF train (Figure 1(B)). Parameter ranges ([begin:step:end]) to build the dictionary were $\text{T}_{\text{1,long}}$ = [20:2:74] ms, $\text{T}_{\text{1,short}}$ = [20:2:74] ms, $\text{T}_{\text{2,long}}^{*}$ = [10:2:66] ms, $\text{T}_{\text{2,short}}^{*}$ = [0.5:0.5:2.0, 2:2:66] ms, $\Delta\text{B}_{\text{1}}^{+}$ factor = [0.7:0.1:1.3] (applied as a multiplying factor to the RF amplitude FA) and $\Delta\text{f}_{\text{0}}$ = [-60:10:60] Hz. For $\text{T}_{\text{1}}$ quantification, we originally assumed that T1,long = T1,short. To insure that spectral densities where $J(0)\neq J(\omega)$ were included in the dictionary, a constant (±$\pm\Delta\text{T}_{\text{1}}$) that added 1 ms to every entry for $\text{T}_{\text{1,long}}$ and subtracted 1 ms to every entry for $\text{T}_{\text{1,short}}$ was included in the simulations. Parameter combinations where $\text{T}_{\text{2,long}}^{*}>\text{T}_{\text{1}}$ and $\text{T}_{\text{2,short}}^{*}>\text{T}_{\text{2,long}}^{*}$ were omitted from the computation. In generating the dictionary, we did not make additional assumptions regarding he ratios between relaxation processes for the central and satellite transitions of the New Aspins $\frac{\text{3}}{\text{2}}$ (i.e. the long and short relaxation components). Neither the dictionary nor the ISTO simulation assume an ideal ratio of 0.6:0.4 for $\text{T}_{\text{2,short}}^{*}$:$\text{T}_{\text{2,long}}^{*}$ due to quadrupolar relaxation processes. We also note here that we are defining the biexponential $\text{T}_{\text{2}}^{*}$ parameters compiled in the dictionary as the effective $\text{T}_{\text{2}}$(35). While $\text{T}_{\text{2}}$ represents decay of the transverse magnetization of the tissue alone, $\text{T}_{\text{2}}^{*}$ includes field inhomogeniety effects(35, 36). In this case, we included $\Delta\text{B}_{\text{1}}^{+}$ factor and $\Delta\text{f}_{\text{0}}$ explicitly in the dictionary to account for the effects of dephasing contributions from field inhomogeneity. Due to the memory expense incurred by the simulation, the dictionary was generated in two parts and concatenated. In total, 831,512 entries were generated. The total simulation time was 6.4 days.
@@
All experiments were performed at 7 T (MAGNETOM, Siemens, Erlangen, Germany) using a 16-channel transmit/receive 1H/23Na RF brain coil constructed in-house (8 1H channels, 8 New Achannels) (37). At the beginning of each scanning session a localizer scan was acquired followed by standard automatic $\text{B}_{\text{0}}$ shimming from the scanner using the 1H channel.
@@
Our test phantom was constructed using a 2.3-L cylinder (outer diameter OD = 20 cm, length = 35 cm) filled with a solution of 70 mM NaCl and which contained seven 50-mL polypropylene cylinders (OD = 30 mm, length = 115 mm) arranged as 6 outer samples each containing a different concentration of NaCl and agar circling the 7th sample placed in the middle and containing a solution of 140 mM NaCl. NaCl concentrations, and thus sodium density ground truth values in the phantoms, were estimated to have a 5% uncertainty due to sample preparation, originating from uncertainty of the scale used to weigh the water, NaCl and Agar before heating and mixing.
A diagram of the phantom with sodium/agar concentrations in the axial orientation is shown in Figure 2. The regions-of-interest (ROI) in the 7 samples were generated from a 3D mask of equal diameter to each sample such that all ROI volumes would be the same.
For New AMRF, one scan consisting of 16 averages with TR = 511 ms, isotropic resolution = 5 mm and isotropic FOV = 320 mm, was acquired with the FLORET parameters: 3 hubs/45 with 100 interleaves/hub, TE = 0.2 ms, readout duration = 5 ms within the composite at the beginning of the pulse train ($\tau_{i}$ = 7.5 ms for $i=1,2$) and 10.230 ms after the next pulses ($\tau_{i}$ = 15 ms for $i\geq 3$), total scan time = 40:52 min.
@@
All images were reconstructed offline in MATLAB. For each channel, raw k-space data acquired during the New AMRF sequence was filtered with a Hamming kernel then reconstructed using gridding (38, 39) to produce a set of 23 complex images with a final nominal resolution 2.5$\times$2.5$\times$2.5 mm3 (128$\times$128$\times$128 matrix). The images from the 8 channels were combined using coil sensitivity profiles as described by Bydder et al. (40). The average time for image reconstruction in the brain across all five volunteers was 42 s, and 38 s in the phantom. For the brain images, an additional denoising step was performed on the complex images using the Marchenko-Pastur method (41, 42, 43). This added another 15 s to the brain image reconstruction time. Supplementary Figure S1 shows 23 axial images for a center slice in the phantom. Supplementary Figures S2 and S3 show 23 reconstructed axial images for a center slice in the brain of volunteer 5.
@@
For $\text{T}_{\text{2}}^{*}$, we applied the biexponential kernel shown in Equation 8 as descibed by Ridley et. al (44, 27).
@@
For each volunteer, images from the 1H MPRAGE DICOM datasets were co-registered to the New AMRF data using SPM12 (UCL, London, UK) (46, 31). The New AMRF image used for co-registration was the one acquired just after the first pulse in the New AMRF train. Tissue probability maps were generated from the normalized and co-registered MPRAGE images and segmented into CSF, GM and WM tissue classes using SPM12 (46, 31). The segmentated regions for CSF, GM and WM were then binarized with a 90% threshold to generate tissue ROI masks in MATLAB. These images are presented in Figure 3. The high threshold was chosen to reduce the likelihood of contamination between different tissues.
@@
Because of the dictionary size, and due to the low SNR of the sodium images, it was possible that more than one match could generate a high correlation for a single voxel. To account for this, we included matches for a subset of the top correlations for each voxel $v$, and generated the final maps by calculating the correlation-weighted parameter $X_{v}$ from the dictionary of values $X_{v,i}$ corresponding to the matching correlation coefficient $w_{v,i}$ according to:
@@
SD in voxel $v$ was calculated as the mean ratio between the mean absolute value of the measured fingerprint in the voxel and the mean absolute value of corresponding fingerprint from the dictionary with correlation coefficient $w_{v,i}$. Once SD was calculated for each voxel $v$ and each correlation coefficient $w_{v,i}$, the resulting SD maps for each $w_{v,i}$, which are now in arbitrary units depending on the data reconstruction algorithm and hardware characterisitics (RF chain, preamplifiers, ADC) were then normalized by dividing them by the highest voxel value in the whole 3D SD dataset. Weighting was then done according to Equation 9, where $X_{v,i}$ was the unweighted SD and $X_{v}$ was the weighted SD. Weighting the SD maps by the correlation coefficients reduces the effect of noise in both the dictionary matching and the normalization processes used to generate the SD maps.